# A Comparison of Galacticus and COZMIC WDM Subhalo Populations

### v1

@@ 1 Introduction
The primary factor that influences the free-streaming scale of warm dark matter is the mass of the warm dark matter particle, $m_{\text{WDM}}$ (Schneider et al., 2013). A wide range of methods have been used to constrain the warm dark matter particle mass, including Lyman-$\alpha$ forest observations (Ballesteros et al., 2021; Garcia-Gallego et al., 2025), dwarf galaxy counts (Dekker et al., 2022), stellar stream features (Banik et al., 2018), strong gravitational lensing (Gilman et al., 2020), and the cosmic reionization history (Tan et al., 2016; Lopez-Honorez et al., 2017). Many of these approaches utilize simulations of WDM to analyze its matter distribution through summary statistics. Different modeling approaches for WDM can lead to subtle differences in the corresponding summary statistics (Avila-Reese et al., 2001; Wang & White, 2007; van den Bosch & Jiang, 2016). Understanding the underlying assumptions and methodologies of each simulation is crucial for interpreting and comparing these statistical discrepancies.
@@
One class of model that is able to generate halo populations more rapidly than $N$-body/hydrodynamical simulations is semi-analytic models (SAMs). SAMs simulate the formation and evolution of dark matter halos (and galaxies) by applying analytic descriptions for certain physical processes to increase the computational efficiency of the model. This computational efficiency is the primary advantage of SAMs, and enables the generation of a large number of realizations—significantly more than would be feasible with full $N$-body simulations—allowing statistical properties to be measured within a reasonable time frame (Henriques et al., 2009; Benson & Bower, 2010; Bower et al., 2010). The main drawback of using SAMs is that there is a greater degree of approximation involved compared to working with N-body simulations. Several studies have shown that SAMs are able to obtain reasonably accurate summary statistics, when compared to N-body simulations (Benson et al., 2002b; Taylor & Babul, 2004; Zentner et al., 2005; Pullen et al., 2014; Yang et al., 2020; Nadler et al., 2023). SAMs have also been used to generate input data for machine learning algorithms, enabling the rapid production of statistically representative halo populations (Kamdar et al., 2016; Elliott et al., 2021; Lonergan et al., 2025).
@@
In this paper, we compare modeling approaches of WDM dark matter subhalo populations made by the semi-analytic model Galacticus (Benson, 2012)  [^1] https://github.com/galacticusorg/galacticus, we use revision 79e9402. and the COsmological ZooM-in simulations with Initial Conditions beyond CDM (COZMIC) N-body simulation suite (Nadler et al., 2025; An et al., 2025). Work has already been done to compare Galacticus CDM halo populations against simulation data (Yang et al., 2020; Nadler et al., 2023), as well as to compare predictions for WDM halo populations from another semi-analytic model, SASHIMI-W (Ando, 2023), with simulations (Ono et al., 2025). The results from Ono et al. (2025) primarily focus on comparing subhalo mass functions (SMFs) between models. In this work, we will expand on these results by comparing additional summary statistics as well as including simulations with a wider range of WDM particle masses in our analysis.
@@ 2.1 COZMIC Simulations
We use data from the COZMIC simulation suite (Nadler et al., 2025), which consists of dark matter only N-body zoom-in simulations of Milky Way (MW) mass host halos for beyond-CDM models. We specifically utilize the results for thermal-relic WDM models, where the linear matter power spectrum is characterized by the warm dark matter particle mass. (Sub-)halo populations were generated for $m_{\text{WDM}}=3,4,5,6,6.5,10$ keV models with three halo realizations per model designed to resimulate two halos from the Milky Way-est simulation suite (Buch et al., 2024) plus one halo from Symphony. Hereafter, these models will be referenced as WDMXX, where $X\in\{3,4,5,6,6.5,10\}$. Initial conditions were generated using MUSIC (Hahn & Abel, 2011), and the simulations were run using GADGET-2 (Springel, 2005) from $z=99$ down to $z=0$. In the highest resolution region of each simulation, the particle mass is $m=4.0\times 10^{5}\mathrm{M}_{\odot}$, equivalent to $8192^{3}$ particles in the original simulation volume from which the resimulations were drawn. Requiring at least 300 particles per halo to ensure a well-converged subhalo mass function, this corresponds to a halo mass resolution of $m_{\text{res}}=1.2\times 10^{8}\mathrm{M}_{\odot}$. To ensure other summary statistics, such as the $V_{\text{max}}$ and $R_{\text{max}}$ distributions, are well-converged, a higher 2,000 particle per halo limit is imposed (Nadler et al., 2025). At this limit, the minimum halo mass becomes $m_{\text{res}}=8.0\times 10^{8}\mathrm{M}_{\odot}$ at the fiducial resolution. To accurately resolve summary statistics beneath this mass threshold, a higher resolution resimulation of COZMIC realization Halo004 is used with a particle mass of $m=5.0\times 10^{4}\mathrm{M}_{\odot}$, enabling convergence down to $1.0\times 10^{8}\mathrm{M}_{\odot}$. WDM halos in the COZMIC simulations are identified using the RockStar phase-space halo finder (Behroozi et al., 2012). Full details of the simulation suite are described in Nadler et al. (2025).
@@ 2.2 Symphony Simulations
The primary aim of this work is to analyze warm dark matter subhalo populations. To assess how well Galacticus performs in matching the COZMIC WDM simulations, we also compare Galacticus CDM subhalo populations to CDM halos from the Symphony simulation suite—this will allow us to determine if Galacticus performs equally well for WDM and CDM. Symphony is a set of cosmological zoom-in simulations of CDM halos in the $10^{11}$$10^{15}\mathrm{M}_{\odot}$ range. We specifically make use of the MW mass host halos, where in the most refined region, the mass and spatial resolution are equivalent to those of a uniform simulation with $8192^{3}$ particles in the simulation box and a simulation particle mass of $m=4.0\times 10^{5}\mathrm{M}_{\odot}$. As with the COZMIC data, a higher resolution resimulation was run for realization Halo004 with particle mass of $5.0\times 10^{4}\mathrm{M}_{\odot}$ to ensure the distributions of halo density profiles are well converged down to the $1.0\times 10^{8}\mathrm{M}_{\odot}$ halo mass scale. Among these MW mass halos, one realization, Halo023, was also resimulated by COZMIC for each WDM mass. A complete description of the simulation setup can be found in Nadler et al. (2023).
@@ 2.4 Galacticus Simulations
In order to compare against the COZMIC, Symphony, and Milky Way-est simulation data, we generate subhalo populations using the Galacticus semi-analytic framework (Benson, 2012). In Galacticus, subhalo populations are generated by constructing a merger tree that identifies all progenitor halos, and their sequence of merging, of a halo at $z=0$. The merger tree is constructed through a Monte Carlo (MC) process using branching rates as predicted from the Extended Press-Schechter (EPS) formalism (Bower, 1991; Lacey & Cole, 1994), but modified to match the results of CDM $N$-body simulations following the algorithm of Parkinson et al. (2008). Once the merger tree is constructed, each halo in the tree is numerically evolved forwards in time, yielding a $z=0$ subhalo population. During the evolution process, orbital properties of each halo, such as their 3D position, velocity, bound mass, and density profile are found by solving a set of differential equations that describe the physics that affects subhalos (Nadler et al., 2023; Lonergan et al., 2025). Specifically, Galacticus accounts for non-linear phenomena such as tidal heating, tidal stripping, and dynamical friction of subhalos as has previously been described and validated for CDM (Pullen et al., 2014; Yang et al., 2020; Benson & Du, 2022).
@@
where $k$ is wavenumber, $\alpha$ corresponds to the suppression scale and $\nu=1.049$ from Vogel & Abazajian (2023). For WDM models, we also switch to using a sharp-$k$ window function to compute $\sigma(M)$ (the root-variance in the density field) following Benson et al. (2013) to avoid the creation of spurious halos below the cut-off scale, and with a normalization parameter $a=2.5$ as advocated by that work. Since we retain a top-hat window function for CDM, to ensure that the halo mass function is unchanged on large scales, we scale [^2] Benson et al. (2013) took a somewhat different approach to resolving this issue. They compute halo mass functions and merger rates by numerically solving the excursion set problem, and so introduced a factor into the excursion set barrier function to correct for the change in $\sigma(M)$ on large scales when using a sharp-$k$ window function. Since, in extended Press-Schechter theory, it is only the ratio of the barrier function to $\sigma^{2}(M)$ that appears in equations governing halo merger rates, our approach is equivalent. $\sigma(M)$ by a factor $0.83$.
@@ 3 Results
The subhalo mass function ratio curves for WDM are largely consistent with a ratio of $1$, given the uncertainty in the N-body results arising from halo-to-halo scatter and the limited number of realizations. There is a trend for the ratio to increase above the equality line for higher subhalo masses, although here the uncertainties become quite large [^3] Note that while the same trend is apparent for all WDM masses, this may simply be a result of the COZMIC simulations using the same set of random phases in their initial conditions, with only the amplitudes of the modes changed.. Importantly, there does not seem to be any systematic variation in the ratio as a function of WDM particle mass, indicating that Galacticus is successfully reproducing the changes in the COZMIC N-body subhalo bound mass functions that arise from WDM physics.
For the CDM case (black line), the uncertainties are much smaller (due to the larger number of CDM N-body realizations), and the ratio falls significantly below 1 for low subhalo masses. A similarly low ratio was found by Nadler et al. (2023) when comparing the ratio of Symphony SHMFs to those from the Caterpillar simulations (Griffen et al., 2016), which is represented by the black dotted curve. The Galacticus subhalo evolution model was calibrated against Caterpillar simulation data (Yang et al., 2020), which explains why it produces a similarly low ratio in comparison with the Symphony and COZMIC data. The CDM case also shows the trend of increasing ratio at high subhalo mass (as was previously reported by Nadler et al. 2023). Given the smaller uncertainties, this trend is significant, suggesting that Galacticus does not accurately capture the physics of these higher mass subhalos. Furthermore, this suggests that the corresponding and very similar trend in the WDM cases is also likely a result of inaccuracies in the Galacticus physics implementation at these higher subhalo masses, and can not simply be explained by poor statistics.
@@
The shaded regions show the $1\sigma$ uncertainties on the mean of the N-body estimates arising from halo-to-halo variance. The uncertainty on the Galacticus results is much smaller as a result of the substantially larger number of realizations available. For WDM, the Galacticus results remain largely enclosed within the shaded $1\sigma$ regions, suggesting agreement with COZMIC data at the level of expected statistical fluctuations.
@@
To compare averaged $V_{\text{max}},R_{\text{max}}$ from Galacticus subhalo populations against those from COZMIC, Figure 7 shows the same averaged $V_{\text{max}},R_{\text{max}}$ values as in Figure 5, but now with curves for both Galacticus and COZMIC models. The figure zooms into subhalos with masses between $10^{8}\mathrm{M}_{\odot}$ to $10^{9}\mathrm{M}_{\odot}$ to focus on subhalos most impacted by WDM physics. As noted by Nadler et al. (2023), a 300 particle threshold in the Symphony $N$-body simulations ensures convergence of the subhalo mass function but not necessarily other summary statistics, particularly those related to the internal structure of subhalos. To robustly capture averaged $V_{\text{max}}$ and $R_{\text{max}}$ profiles, a stricter 2,000-particle cut is adopted. The fiducial resolution COZMIC data is plotted above its 2,000 particle mass limit of $m=8.0\times 10^{8}\mathrm{M}_{\odot}$. Below this mass scale, the higher resolution data is shown down to its 2,000 particle threshold.
@@
In Figure 6, the WDM1 model predictably experiences the most significant suppression, resulting in lower collapse redshifts across all mass scales. However, if we examine the degree of suppression as a function of halo mass relative to the other models, some interesting behavior can be discerned. For example, the WDM1 collapse redshift curve is significantly lower than all other models in the $10^{9},10^{10}]\mathrm{M}_{\odot}$ mass range; however, at the lowest masses shown, the WDM1 model is suppressed by an amount very similar to WDM2 and even WDM3 models. As noted above, the collapse redshift is constrained by the infall redshift curves (dotted lines), as halos must satisfy $z_{\text{collapse}}>z_{\text{infall}}$. Infall redshifts correspond to when a halo first merged into a larger system and became a subhalo, and so inherently depend on the merger rates used to construct our merger trees. The merger rate in the extended Press-Schechter theory [^4] Our actual merger tree calculations include the modifier term introduced by Parkinson et al. (2008) to better match merger rates in N-body simulations. We do not include that here for simplicity as it does not qualitatively affect the argument that we present. is given by Lacey & Cole (1993):
@@
For example, considering masses $M_{1}\sim 10^{8}\mathrm{M}_{\odot}$ and $M_{2}\sim 10^{11}$$10^{12}\mathrm{M}_{\odot}$, it can be seen that that, for WDM1, $\sigma_{1}$ and $\sigma_{2}$ are much closer to each other than for CDM (or for most of the other WDM models). Significantly, at higher redshifts, $M_{2}(z)$ decreases, making the difference between $\sigma_{1}$ and $\sigma_{2}$ smaller (and so the term in the merger rate becomes larger). Thus, for extreme WDM models, such as WDM1, merger rates are relatively higher at higher redshifts compared to $z=0$, resulting in a bias toward higher infall redshifts. This forces a similar increase in collapse redshifts for low mass halos, which in turn leads to $R_{\mathrm{max}}$ trending back toward the CDM expectation for these extreme models. The same effect, albeit weaker, is present for WDM3 and causes the downturn in the average $R_{\mathrm{max}}$ at the lowest halo masses shown in Figure 4.

### v2

@@ 1 Introduction
The primary factor that influences the free-streaming scale of warm dark matter is the mass of the warm dark matter particle, $m_{\text{WDM}}$ (Schneider et al., 2013). A wide range of methods have been used to constrain the warm dark matter particle mass, including Lyman-$\alpha$ forest observations (Ballesteros et al., 2021; Garcia-Gallego et al., 2025), dwarf galaxy counts (Dekker et al., 2022), stellar stream features (Banik et al., 2018), strong gravitational lensing (Gilman et al., 2020), and the cosmic reionization history (Tan et al., 2016; Lopez-Honorez et al., 2017). Many of these approaches utilize simulations of WDM to analyze its matter distribution through summary statistics. Different modeling approaches for WDM can lead to subtle differences in the corresponding summary statistics (Avila-Reese et al., 2001; Wang and White, 2007; van den Bosch and Jiang, 2016). Additionally, the WDM particle mass is one of multiple factors which influence the decoupling mechanism of the dark matter particle. A detailed discussion on these phenomenological differences, and how they are implemented in current simulations can be found in Paduroiu (2022). Understanding the underlying assumptions and methodologies of each simulation is crucial for interpreting and comparing these statistical discrepancies (Nierenberg et al., 2016).
@@
One class of model that is able to generate halo populations more rapidly than $N$-body/hydrodynamical simulations is semi-analytic models (SAMs). SAMs simulate the formation and evolution of dark matter halos (and galaxies) by applying analytic descriptions for certain physical processes to increase the computational efficiency of the model. This computational efficiency is the primary advantage of SAMs, and enables the generation of a large number of realizations—significantly more than would be feasible with full $N$-body simulations—allowing statistical properties to be measured within a reasonable time frame (Henriques et al., 2009; Benson and Bower, 2010; Bower et al., 2010). The main drawback of using SAMs is that there is a greater degree of approximation involved compared to working with N-body simulations. Several studies have shown that SAMs are able to obtain reasonably accurate summary statistics, when compared to N-body simulations (Benson et al., 2002b; Taylor and Babul, 2004; Zentner et al., 2005; Pullen et al., 2014b; Yang et al., 2020b; Nadler et al., 2023b). SAMs have also been used to generate input data for machine learning algorithms, enabling the rapid production of statistically representative halo populations (Kamdar et al., 2016; Elliott et al., 2021; Lonergan et al., 2025).
@@
In this paper, we compare modeling approaches of WDM dark matter subhalo populations made by the semi-analytic model Galacticus (Benson, 2012)  [^1] https://github.com/galacticusorg/galacticus, we use revision 79e9402. and the COsmological ZooM-in simulations with Initial Conditions beyond CDM (COZMIC) N-body simulation suite (Nadler et al., 2025; An et al., 2025). Work has already been done to compare Galacticus CDM halo populations against simulation data (Yang et al., 2020a; Nadler et al., 2023a), as well as to compare predictions for WDM halo populations from another semi-analytic model, SASHIMI-W (Ando, 2023), with simulations (Ono et al., 2025). The results from Ono et al. (2025) primarily focus on comparing subhalo mass functions (SMFs) between models. In this work, we will expand on these results by comparing additional summary statistics as well as including simulations with a wider range of WDM particle masses in our analysis.
@@ 2.1 COZMIC Simulations
We use data from the COZMIC simulation suite (Nadler et al., 2025), which consists of dark matter only N-body zoom-in simulations of Milky Way (MW) mass host halos for beyond-CDM models. We specifically utilize the results for thermal-relic WDM models, where the linear matter power spectrum is characterized by the warm dark matter particle mass. (Sub-)halo populations were generated for $m_{\text{WDM}}=3,4,5,6,6.5,10$ keV models with three halo realizations per model designed to resimulate two halos from the Milky Way-est simulation suite (Buch et al., 2024) plus one halo from Symphony. Hereafter, these models will be referenced as WDMXX, where $X\in\{3,4,5,6,6.5,10\}$. Initial conditions were generated using MUSIC (Hahn and Abel, 2011), and the simulations were run using GADGET-2 (Springel, 2005) from $z=99$ down to $z=0$. In the highest resolution region of each simulation, the particle mass is $m=4.0\times 10^{5}\mathrm{M}_{\odot}$, equivalent to $8192^{3}$ particles in the original simulation volume from which the resimulations were drawn. Requiring at least 300 particles per halo to ensure a well-converged subhalo mass function, this corresponds to a halo mass resolution of $m_{\text{res}}=1.2\times 10^{8}\mathrm{M}_{\odot}$. To ensure other summary statistics, such as the $V_{\text{max}}$ and $R_{\text{max}}$ distributions, are well-converged, a higher 2,000 particle per halo limit is imposed (Nadler et al., 2025). At this limit, the minimum halo mass becomes $m_{\text{res}}=8.0\times 10^{8}\mathrm{M}_{\odot}$ at the fiducial resolution. To accurately resolve summary statistics beneath this mass threshold, a higher resolution resimulation of COZMIC realization Halo004 is used with a particle mass of $m=5.0\times 10^{4}\mathrm{M}_{\odot}$, enabling convergence down to $1.0\times 10^{8}\mathrm{M}_{\odot}$. WDM halos in the COZMIC simulations are identified using the RockStar phase-space halo finder (Behroozi et al., 2012). Full details of the simulation suite are described in Nadler et al. (2025).
@@ 2.2 Symphony Simulations
The primary aim of this work is to analyze warm dark matter subhalo populations. To assess how well Galacticus performs in matching the COZMIC WDM simulations, we also compare Galacticus CDM subhalo populations to CDM halos from the Symphony simulation suite—this will allow us to determine if Galacticus performs equally well for WDM and CDM. Symphony is a set of cosmological zoom-in simulations of CDM halos in the $10^{11}$$10^{15}\mathrm{M}_{\odot}$ range. We specifically make use of the MW mass host halos, where in the most refined region, the mass and spatial resolution are equivalent to those of a uniform simulation with $8192^{3}$ particles in the simulation box and a simulation particle mass of $m=4.0\times 10^{5}\mathrm{M}_{\odot}$. As with the COZMIC data, a higher resolution resimulation was run for realization Halo004 with particle mass of $5.0\times 10^{4}\mathrm{M}_{\odot}$ to ensure the distributions of halo density profiles are well converged down to the $1.0\times 10^{8}\mathrm{M}_{\odot}$ halo mass scale. Among these MW mass halos, one realization, Halo023, was also resimulated by COZMIC for each WDM mass. A complete description of the simulation setup can be found in Nadler et al. (2023a).
@@ 2.4 Galacticus Simulations
In order to compare against the COZMIC, Symphony, and Milky Way-est simulation data, we generate subhalo populations using the Galacticus semi-analytic framework (Benson, 2012). In Galacticus, subhalo populations are generated by constructing a merger tree that identifies all progenitor halos, and their sequence of merging, of a halo at $z=0$. The merger tree is constructed through a Monte Carlo (MC) process using branching rates as predicted from the Extended Press-Schechter (EPS) formalism (Bower, 1991; Lacey and Cole, 1994), but modified to match the results of CDM $N$-body simulations following the algorithm of Parkinson et al. (2008). Once the merger tree is constructed, each halo in the tree is numerically evolved forwards in time, yielding a $z=0$ subhalo population. During the evolution process, orbital properties of each halo, such as their 3D position, velocity, bound mass, and density profile are found by solving a set of differential equations that describe the physics that affects subhalos (Nadler et al., 2023a; Lonergan et al., 2025). Specifically, Galacticus accounts for non-linear phenomena such as tidal heating, tidal stripping, and dynamical friction of subhalos as has previously been described and validated for CDM (Pullen et al., 2014a; Yang et al., 2020a; Benson and Du, 2022). It should be noted that while Galacticus uses a bottom up model for the structure formation of WDM halos, other studies point out that the structure formation can be more complex depending on the redshift and mass scale at which a halo forms (Paduroiu, 2022). It would be interesting to consider structure formation histories alternative to the “traditional” merger tree scenario in future work.
@@
where $k$ is wavenumber, $\alpha$ corresponds to the suppression scale and $\nu=1.049$ from Vogel and Abazajian (2023). For WDM models, we also switch to using a sharp-$k$ window function to compute $\sigma(M)$ (the root-variance in the density field) following Benson et al. (2013) to avoid the creation of spurious halos below the cut-off scale, and with a normalization parameter $a=2.5$ as advocated by that work. Since we retain a top-hat window function for CDM, to ensure that the halo mass function is unchanged on large scales, we scale [^2] Benson et al. (2013) took a somewhat different approach to resolving this issue. They compute halo mass functions and merger rates by numerically solving the excursion set problem, and so introduced a factor into the excursion set barrier function to correct for the change in $\sigma(M)$ on large scales when using a sharp-$k$ window function. Since, in extended Press-Schechter theory, it is only the ratio of the barrier function to $\sigma^{2}(M)$ that appears in equations governing halo merger rates, our approach is equivalent. $\sigma(M)$ by a factor of $0.83$.
@@ 3 Results
The subhalo mass function ratio curves for WDM are largely consistent with a ratio of $1$, given the uncertainty in the N-body results arising from halo-to-halo scatter and the limited number of realizations. There is a trend for the ratio to increase above the equality line for higher subhalo masses, although here the uncertainties become quite large. Importantly, there does not seem to be any systematic variation in the ratio as a function of WDM particle mass, which may simply be a result of the COZMIC simulations using the same set of random phases in their initial conditions, with only the amplitudes of the modes changed. The lack of $m_{\text{WDM}}$ dependence in the SMF ratio curves indicates that Galacticus is successfully reproducing the changes in the COZMIC N-body subhalo bound mass functions that arise from WDM physics.
For the CDM case (black line), the uncertainties are much smaller (due to the larger number of CDM N-body realizations), and the ratio falls significantly below 1 for low subhalo masses. A similarly low ratio was found by Nadler et al. (2023a) when comparing the ratio of Symphony SHMFs to those from the Caterpillar simulations (Griffen et al., 2016), which is represented by the black dotted curve. The Galacticus subhalo evolution model was calibrated against Caterpillar simulation data (Yang et al., 2020a), which explains why it produces a similarly low ratio in comparison with the Symphony and COZMIC data. The CDM case also shows the trend of increasing ratio at high subhalo mass (as was previously reported by Nadler et al. 2023a). Given the smaller uncertainties, this trend is significant, suggesting that Galacticus does not accurately capture the physics of these higher mass subhalos. Furthermore, this suggests that the corresponding and very similar trend in the WDM cases is also likely a result of inaccuracies in the Galacticus physics implementation at these higher subhalo masses, and can not simply be explained by poor statistics.
@@
The shaded regions show the $1\sigma$ uncertainties on the mean of the N-body estimates arising from halo-to-halo variance. The uncertainty on the Galacticus results is much smaller as a result of the substantially larger number of realizations available. For WDM, the Galacticus results remain largely enclosed within the shaded $1\sigma$ regions, suggesting agreement with COZMIC data at the level of expected statistical fluctuations. It should be noted that even at low radii, both the Galacticus and COZMIC radial distribution curves do not scale monotonically with WDM particle mass. The results for the Galacticus curves are statistically robust, as the difference in WDM models is larger than the statistical uncertainty around a given model. Details around this non-monotonic behavior will be discussed further in section 4.
@@
To compare averaged $V_{\text{max}},R_{\text{max}}$ from Galacticus subhalo populations against those from COZMIC, Figure 6 shows the same averaged $V_{\text{max}},R_{\text{max}}$ values as in Figure 5, but now with curves for both Galacticus and COZMIC models. The figure zooms into subhalos with masses between $10^{8}\mathrm{M}_{\odot}$ to $10^{9}\mathrm{M}_{\odot}$ to focus on subhalos most impacted by WDM physics. As noted by Nadler et al. (2023a), a 300 particle threshold in the Symphony $N$-body simulations ensures convergence of the subhalo mass function but not necessarily other summary statistics, particularly those related to the internal structure of subhalos. To robustly capture averaged $V_{\text{max}}$ and $R_{\text{max}}$ profiles, a stricter 2,000-particle cut is adopted. The fiducial resolution COZMIC data is plotted above its 2,000 particle mass limit of $m=8.0\times 10^{8}\mathrm{M}_{\odot}$. Below this mass scale, the higher resolution data is shown down to its 2,000 particle threshold.
@@
In Figure 7, the WDM1 model predictably experiences the most significant suppression, resulting in lower collapse redshifts across all mass scales. However, if we examine the degree of suppression as a function of halo mass relative to the other models, some interesting behavior can be discerned. For example, the WDM1 collapse redshift curve is significantly lower than all other models in the $10^{9},10^{10}]\mathrm{M}_{\odot}$ mass range; however, at the lowest masses shown, the WDM1 model is suppressed by an amount very similar to WDM2 and even WDM3 models. As noted above, the collapse redshift is constrained by the infall redshift curves (dotted lines), as halos must satisfy $z_{\text{collapse}}>z_{\text{infall}}$. Infall redshifts correspond to when a halo first merged into a larger system and became a subhalo, and so inherently depend on the merger rates used to construct our merger trees. The merger rate in the extended Press-Schechter theory [^3] Our actual merger tree calculations include the modifier term introduced by Parkinson et al. (2008) to better match merger rates in N-body simulations. We do not include that here for simplicity as it does not qualitatively affect the argument that we present. is given by Lacey and Cole (1993):
@@
For example, considering masses $M_{1}\sim 10^{8}\mathrm{M}_{\odot}$ and $M_{2}\sim 10^{11}$$10^{12}\mathrm{M}_{\odot}$, it can be seen that, for WDM1, $\sigma_{1}$ and $\sigma_{2}$ are much closer to each other than for CDM (or for most of the other WDM models). Significantly, at higher redshifts, $M_{2}(z)$ decreases, making the difference between $\sigma_{1}$ and $\sigma_{2}$ smaller (and so the term in the merger rate becomes larger). Thus, for extreme WDM models, such as WDM1, merger rates are relatively higher at higher redshifts compared to $z=0$, resulting in a bias toward higher infall redshifts. This forces a similar increase in collapse redshifts for low mass halos, which in turn leads to $R_{\mathrm{max}}$ trending back toward the CDM expectation for these extreme models. The same effect, albeit weaker, is present for WDM3 and causes the downturn in the average $R_{\mathrm{max}}$ at the lowest halo masses shown in Figure 4.